OpenStudy (anonymous):

Anne Katz, owner of Katz Sport Shop, loans $8,000 to Shelley Slater to help her open an art shop. Shelley plans to repay Anne at the end of 8 years with interest compounded semiannually at 8 percent. At the end of 8 years, Anne will receive: (Use the tables in the handbook)

OpenStudy (anonymous):

8000*(1+.08/2)^(8*2) =8000*(1.04)^16 =$14,983.85